# CRESSim–MPM: A Material Point Method Library for Surgical Soft Body Simulation with Cutting and Suturing

### v2

@@ Abstract
A number of recent studies have focused on developing surgical simulation platforms to train machine learning (ML) agents or models with synthetic data for surgical assistance. While existing platforms excel at tasks such as rigid body manipulation and soft body deformation, they struggle to simulate more complex soft body behaviors like cutting and suturing. A key challenge lies in modeling soft body fracture and splitting using the finite-element method (FEM), which is the predominant approach in current platforms. Additionally, the two-way suture needle/thread contact inside a soft body is further complicated when using FEM. In this work, we use the material point method (MPM) for such challenging simulations and propose new rigid geometries and soft-rigid contact methods specifically designed for them. We introduce CRESSim-MPM, a GPU-accelerated MPM library that integrates multiple MPM solvers and incorporates surgical geometries for cutting and suturing, serving as a specialized physics engine for surgical applications. It is further integrated into Unity, requiring minimal modifications to existing projects for soft body simulation. We demonstrate the simulator’s capabilities in real-time simulation of cutting and suturing on soft tissue and provide an initial performance evaluation of different MPM solvers when simulating varying numbers of particles.
@@ I INTRODUCTION
While there are numerous high-performance and real-time simulators for other robotics applications such as autonomous driving and general robot manipulation, there are limited resources in terms of surgical robotics simulation. Although simulating rigid surgical instruments is usually not challenging, the simulation of surgical scenes and operations requires special simulation capabilities that are less present in other robotics fields.
@@
To the best of our knowledge, this is also the first work to simulate soft tissue suturing using the MPM. Our method naturally couples 1D suturing geometries (i.e., an arc-shaped needle with segmented suture thread) with soft tissue dynamics, enabling two-way coupling between MPM soft bodies and 1D rigid bodies. This formulation also lays the groundwork for incorporating other solvers such as position-based dynamics (PBD) for future extensions. The use of MPM further allows for easy future integration of fluids, such as blood and smoke in surgeries, as MPM naturally supports fluid simulation.
CRESSim-MPM is part of the CRESSim project [^1] https://tbs-ualberta.github.io/CRESSim/ [12, 6] and will be openly available to facilitate surgical robotics and surgical skills training research.
@@ II-A Surgical Simulators and Surgical Robotics
Surgical simulator development has mainly been influenced by two streams. The first one originates from the need for surgical training and skills assessment within a simulated environment, with the users being surgeons, residents, and medical researchers. Such simulators are typically capable of simulating a wide range of surgical training tasks realistically, such as cutting, suturing, and cauterization. Examples of these are the commercial ones such as da Vinci SimNow [^2] https://www.intuitive.com/products-and-services/da-vinci/learning/simnow/ and the open-source ones such as iMSTK [^3] https://www.imstk.org/. However, they are less used by robotics researchers due to their proprietary nature, the lack of real-time performance, or difficulty in integrating with existing software, such as the Robot Operating System (ROS).
The second stream arises from the recent trend in surgical robotics research, particularly for surgical automation and autonomy, which relies on simulators for benchmarking, synthesizing data, and trial-and-error learning. Existing studies such as AMBF [13], ATAR [14], and V-Rep Simulator for the dVRK [15] provide diverse applications for surgical robotics research. Other work such as ORBIT-Surgical [16], Surgical Gym [17], FF-SRL [18], LapGym [19], AMBF-RL [20], SurRoL [21], dVRL [22], and UnityFlexML [1] has specific focus on machine learning applications. These platforms typically rely on an existing physics engine, such as PhysX (used in Nvidia Omniverse), FleX, SOFA, and Bullet. Therefore, their simulation capability largely depends on the physics solvers. For instance, Bullet is primarily a rigid body engine and only provides limited FEM soft body support. PhysX 5 supports FEM, but interactions such as cutting and suturing in soft bodies have not been realistically achieved, possibly due to the limitations of FEM. A Recent study [23] integrates basic MPM support into SurRoL using MLS-MPM but does not address cutting and suturing contact methods. In contrast, our work focuses on developing a comprehensive engine-level MPM library specifically designed for simulating surgical soft body manipulations.
@@ II-B Real-Time Soft Body Simulation
PBD [24] and extended PBD (XPBD) [25] are efficient and numerically stable alternatives, which solve only positional constraints without relying on force integration, making the simulation highly stable. However, this also means less physically accurate behavior and the realism of deformations depends on the formulation of constraints rather than the underlying physics. Although this is acceptable for computer animation, the application in biomedical and robotics research poses additional challenges, as careful parameter tuning is needed to match the simulation with real-world tissue behavior, as in [26]. Additionally, forming positional constraints can sometimes be challenging for certain types of materials, such as viscoelastic or elastoplastic models, which is common in biomedical applications.
MPM overcomes many disadvantages of FEM while preserving accuracy and versatility for simulating a wide range of material behaviors. MPM is considered within the particle-in-cell (PIC) methods family, where particles carry physical quantities and a background grid is used for computations. The initial PIC method [27] was developed for fluid mechanics, with later improvements by the fluid implicit particle method (FLIP) [28]. MPM extends the methods to simulating solids and multiphase materials, including elastic and hyperelastic materials, granular materials (e.g., snow and sand), and elastoplastic materials such as biological tissues. It is particularly famous for snow simulation in Disney’s film Frozen [29]. In addition, since MPM does not explicitly use a mesh, topological changes are not a concern, making it easier to simulate cutting and fracture. The MPM and the PIC methods are closely related and have numerous variations such as generalized interpolation material point (GIMP) [30], convected particle domain interpolation (CPDI) [31], affine particle-in-cell (APIC) [32], moving least squares MPM (MLS-MPM) [33], and more recently position-based MPM (PB-MPM) [34]. However, the method is less investigated and adopted for surgical simulation, possibly due to its recent emergence and the lack of a high-performance library that supports a range of MPM formulations.
@@ III-A Standard MPM
where $\mathbf{f}_{I}^{int}$ is the internal force and $\mathbf{f}_{I}^{ext}$ is the external force. $V_{i}$ is the particle volume, $\boldsymbol{\sigma}_{i}$ is the stress tensor, $\mathbf{b}$ is the body force. $\mathbf{b}=\mathbf{g}$ when only gravity is applied. $\nabla\phi$ is the gradient of the weighting function.
@@
Equation (8) is specific to the Neo-Hookean model for calculating the Cauchy stress, where $\mu$ and $\lambda$ are the Lamé parameters. Some methods use the Kirchhoff stress instead.
One benefit of MPM is that contact between multiple soft bodies is naturally handled, as each body is represented by a cluster of particles interacting through the shared background grid. This eliminates the need for explicit collision detection.
Since various improved algorithms are now more commonly used, standard MPM is implemented in the library only as a reference and does not support all features.
@@ III-B Moving Least Squares MPM
The moving least squares MPM (MLS-MPM) uses APIC-style transfers and unifies it with force computation by using a moving least squares shape function. This leads to modifications of the velocity and force transfers in the standard MPM formulation. For the internal force,
@@ III-C Position-Based MPM
PB-MPM, proposed in 2024 [34], is a relatively new method that formulates MPM as a constraint-solving problem similar to PBD, allowing larger integration steps and better numerical stability. The modifications needed to convert any MPM algorithm to PB-MPM are minimal: at each simulation step, multiple P2G-G2P cycles are conducted to iteratively solve a candidate particle velocity gradient. In P2G, instead of calculating the stress, a material constraint is solved for all particles independently. For example, for an elastic material with MLS-MPM formulation, the co-rotational constraint-solving step includes
@@ III-D Computational Speed Comparison
We provide an initial performance evaluation with different numbers of particles for the implemented solvers on Nvidia RTX 4090 (compute capability 8.9), with CUDA Toolkit version 12.4.1. For all solvers, each step advances the simulation by $0.02$ second. For MLS-MPM and standard MPM, each step contains 10 integration sub-steps with $\Delta t=0.002$. For PB-MPM, each step uses $\Delta t=0.02$ with 10 or 20 iterations. The grid size is fixed in all situations with $175,616$ nodes. The results are shown in Fig. 3. While comparing with existing implementations is reasonable, this work does not aim to outperform existing code in computational speed, as our main focus is on including multiple MPM solver options within one library. Further code optimization and performance benchmarking are planned as future work.
@@ IV-A Soft-Rigid Coupling
Figure 4: Grid level collision with SDF. Nodal velocities are modified according to their location inside a rigid body and the contact model.
Two-way coupling between MPM soft bodies and externally-solved rigid bodies can be achieved at the grid level [29, 33]. The most straightforward approach is to use a signed distance field (SDF) for each rigid body and modify the predicted nodal velocity based on the contact model if a node is inside the rigid geometry. This is similar to a predictor-corrector method [10], as shown in Fig. 4. We set the nodal velocities to zero along the normal direction, and apply frictional correction to the tangential component, unless the node is separating from the rigid body. Mathematically, the relative velocity between a node and a rigid body is $\mathbf{v}_{rel}=\mathbf{v}_{I}-\mathbf{v}_{r}$, with $\mathbf{v}_{r}$ being the rigid body’s velocity. Given a rigid surface normal $\mathbf{n}$, $\mathbf{v}_{rel}$ is decomposed into the tangent component $\mathbf{v}_{tg}$ and the normal component $v_{n}$:
@@
Figure 5: Soft body coupling with rigid bodies of different masses.
@@ IV-B Geometries and Contact Method for Cutting
Additionally, since a knife or scalpel can cut with only the edge side instead of the spine (blunt) side, it is inappropriate to have a geometry that cuts in all directions. Considering this, we apply a circular-shaped distance field to the spine side and use sticky contact (zero relative velocity) if a node is in this region. When the spine side touches an MPM soft body, it generates contact similar to a cylinder colliding with the soft body, as shown in Fig. 6. The SDF of its cross-section is illustrated in Fig. 7. The geometry is named slicer in the library, which includes QuadSlicer and TriangleMeshSlicer.
@@
Figure 7: SDF for a slicer geometry.
@@ IV-C Geometries and Contact Method for Suturing
In this work, we assume that the suture thread can be discretized into connected straight-line segments. To achieve suturing, we introduce the arc and the line segment geometries, which can model an arc-shaped suture needle and a discrete thread segment. They are both 1D geometries with SDFs. The signed distance is calculated analytically from a given query point to its nearest projected point on the geometry, which is always positive. The surface normal $\mathbf{n}$ is the normalized vector from the projected point to the query point. Nodal velocities inside the SDF region are corrected differently than for regular geometries. The relative velocity is decomposed into three parts: one along the arc or line tangent, one that lies on the cross-section of the geometry, and one along the surface normal, as shown in Fig. 8, and only the tangent component is kept. Mathematically, the relative velocity $\mathbf{v}_{rel}$ is decomposed to
@@
where $\mathbf{v}_{tg1}$ is the true tangent velocity along the arc or line. Nodal velocity correction is applied by replacing $\mathbf{v}_{tg}$ with $\mathbf{v}_{tg1}$ in (18). This ensures that the rigid geometry can travel freely inside a soft body if the motion is along the geometry (e.g., inserting a needle) while preventing lateral motion, as shown in Fig. 9. Geometries for suturing include Arc and ConnectedLineSegments, where the latter is a generalized geometry of a line segment.
@@ V-A Object Model and Simulation Lifecycle
The library follows an object model similar to the one in PhysX, such that integration with other PhysX-based software, including Unity, is trivial, as shown in Fig. 10.
@@ V-B Implementation Details and Usage
The library is implemented in CUDA C++ and is dependent only on the standard C++ library and the CUDA runtime. 3 types of Solver discussed in Section III are currently included. ParticleObject represents an object of material points simulated by MPM, and Shape is a rigid body. SimulationFactory is used to create all objects. C++ Users should only include the header simulation_factory.cuh. A backend C-style API engine is also provided to allow straightforward integration into existing software.
@@
Figure 11: Soft tissue cutting with a scalpel moving right then down.
@@ V-C Unity Integration
Integration into Unity is straightforward using the C-style APIs. Only a few modifications are needed to allow the coupling between the MPM soft body and rigid bodies simulated in Unity’s default engine. For each rigid body in Unity, a corresponding CRESSim-MPM Shape is created for two-way coupling with the soft body. During each step, a rigid body is simulated first with the default engine, after which the corresponding Shape’s velocity is updated. The MPM engine is then advanced, recording a total momentum change of the contact grids with the rigid body. This is then used to apply a momentum impulse on the rigid body in Unity. A Unity plugin is provided as part of CRESSim-MPM.

### v3

@@ Abstract
A number of recent studies have focused on developing surgical simulation platforms to train machine learning (ML) agents or models with synthetic data for surgical assistance. While existing platforms excel at tasks such as rigid body manipulation and soft body deformation, they struggle to simulate more complex soft body behaviors like cutting and suturing. A key challenge lies in modeling soft body fracture and splitting using the finite-element method (FEM), which is the predominant approach in current platforms. Additionally, the two-way suture needle/thread contact inside a soft body is further complicated when using FEM. In this work, we use the material point method (MPM) for such challenging simulations and propose new rigid geometries and soft-rigid contact methods specifically designed for them. We introduce CRESSim-MPM, a GPU-accelerated MPM library that integrates multiple MPM solvers and incorporates surgical geometries for cutting and suturing, serving as a specialized physics engine for surgical applications. It is further integrated into Unity, requiring minimal modifications to existing projects for soft body simulation. We demonstrate the simulator’s capabilities in real-time simulation of cutting and suturing on soft tissue and provide an initial performance evaluation of different MPM solvers when simulating varying numbers of particles. The source code is available at https://github.com/yafei-ou/CRESSim-MPM.
@@ I INTRODUCTION
While there are numerous high-performance and real-time simulators for other robotics applications, such as autonomous driving (e.g., CARLA) and general robot manipulation (e.g., Isaac Sim), there are limited resources in terms of surgical robotics simulation. Although simulating rigid surgical instruments is usually not challenging, the simulation of surgical scenes and operations requires specific simulation capabilities less present in other robotics fields.
@@
To the best of our knowledge, this is also the first work to simulate soft tissue suturing using the MPM. Our method couples 1D suturing geometries (i.e., an arc-shaped needle with segmented suture thread) with soft tissue dynamics, enabling two-way coupling between MPM soft bodies and 1D rigid bodies. This formulation also lays the groundwork for incorporating other solvers, such as position-based dynamics (PBD), for future extensions. The use of MPM further allows for easy future integration of fluids, such as blood and smoke in surgeries, as MPM naturally supports fluid simulation [10].
CRESSim-MPM is part of the CRESSim project [^1] https://tbs-ualberta.github.io/CRESSim/ [12, 6].
@@ II-A Surgical Simulators and Surgical Robotics
Surgical simulator development, particularly in the area of software-based (virtual) simulators, has mainly been influenced by two streams. The first one originates from the need for surgical training and skills assessment within a simulated environment, with the users being surgeons, residents, and medical researchers. This category primarily includes virtual simulators and hybrid simulators that are capable of simulating a wide range of surgical training tasks, such as cutting, suturing, and cauterization. Examples of these are the commercial ones, such as da Vinci SimNow [^2] https://www.intuitive.com/products-and-services/da-vinci/learning/simnow/ and the open-source ones such as iMSTK [^3] https://www.imstk.org/. However, they are less used by robotics researchers due to their proprietary nature, the lack of real-time performance, or difficulty in integrating with existing software, such as the Robot Operating System (ROS).
The second stream arises from the recent trend in surgical robotics research, particularly for surgical automation and autonomy. Studies including AMBF [13], ATAR [14], and V-Rep Simulator for the dVRK [15] provide diverse applications for surgical robotics research, such as for benchmarking, synthesizing data, and trial-and-error learning. Other works such as ORBIT-Surgical [16], Surgical Gym [17], FF-SRL [18], LapGym [19], AMBF-RL [20], SurRoL [21], dVRL [22], and UnityFlexML [1] has specific focus on machine learning applications. These platforms typically rely on an existing physics engine, such as PhysX (used in Nvidia Omniverse), FleX, SOFA, and Bullet. Therefore, their simulation capability largely depends on the physics solvers. For instance, Bullet is primarily a rigid body engine and only provides limited FEM soft body support. Both PhysX 5 and SOFA use FEM for soft body simulation. While PhysX 5 supports FEM, it has yet to achieve realistic interactions such as cutting and suturing. SOFA does support cutting but incurs high computational costs. These challenges may stem from inherent limitations of FEM itself, as discussed earlier. A recent study [23] integrates basic MPM support into SurRoL using MLS-MPM but does not address cutting and suturing contact methods. In contrast, our work focuses on developing a comprehensive engine-level MPM library specifically designed for simulating surgical soft-body manipulations.
@@ II-B Real-Time Soft Body Simulation
PBD [24] and extended PBD (XPBD) [25] are efficient and numerically stable alternatives, which solve only positional constraints without relying on force integration, making the simulation highly stable. However, this also means less physically accurate behavior, and the realism of deformations depends on the formulation of constraints rather than the underlying physics. Although this is acceptable for computer animation, the application in biomedical and robotics research poses additional challenges, as careful parameter tuning is needed to match the simulation with real-world tissue behavior, as in [26]. Additionally, forming positional constraints can sometimes be challenging for certain types of materials, such as viscoelastic or elastoplastic models, which are common in biomedical applications.
MPM overcomes many disadvantages of FEM while preserving accuracy and versatility for simulating a wide range of material behaviors. MPM is considered within the particle-in-cell (PIC) methods family, where particles carry physical quantities and a background grid is used for computations. The initial PIC method [27] was developed for fluid mechanics, with later improvements by the fluid implicit particle method (FLIP) [28]. MPM extends the methods to simulating solids and multiphase materials, including elastic and hyperelastic materials, granular materials (e.g., snow–used by Disney in their film Frozen [29]–and sand), and elastoplastic materials such as biological tissues. In addition, since MPM does not explicitly use a mesh, topological changes are not a concern, making it easier to simulate cutting and fracture. The MPM and the PIC methods are related and have numerous variations such as generalized interpolation material point (GIMP) [30], convected particle domain interpolation (CPDI) [31], affine particle-in-cell (APIC) [32], moving least squares MPM (MLS-MPM) [33], and more recently position-based MPM (PB-MPM) [34]. However, MPM is less investigated and adopted for surgical simulation, possibly due to the lack of a high-performance library specifically built for surgical applications.
@@ III-A Standard MPM
where $\mathbf{f}_{I}^{int}$ is the internal force and $\mathbf{f}_{I}^{ext}$ is the external force. $V_{i}$ is the particle volume, $\boldsymbol{\sigma}_{i}$ is the stress tensor, $\mathbf{b}$ is the body force. $\mathbf{b}=\mathbf{g}$ when only gravity is applied. $\nabla\phi$ is the gradient of the shape function.
@@
$\mathbf{I}$ denotes the identity matrix. Equation (8) is specific to the Neo-Hookean model for calculating the Cauchy stress, where $\mu$ and $\lambda$ are the Lamé parameters. Some implementations use the Kirchhoff stress instead.
One benefit of MPM is that contact between multiple soft bodies is inherently handled, as each body is represented by a cluster of particles interacting through the shared background grid. This eliminates the need for explicit collision detection.
Since more advanced algorithms are now commonly used, standard MPM is implemented in this work only as a reference with limited feature support.
@@ III-B Moving Least Squares MPM
The moving least squares MPM (MLS-MPM) [33] uses APIC-style transfers and unifies it with force computation by using a moving least squares shape function. This leads to modifications of the velocity and force transfers in the standard MPM formulation. For the internal force,
@@ III-C Position-Based MPM
PB-MPM, proposed in 2024 [34], is a new method that formulates MPM as a constraint-solving problem similar to PBD, allowing larger integration steps and better numerical stability. The modifications needed to convert any MPM algorithm to PB-MPM are minimal: at each simulation step, multiple P2G-G2P cycles are conducted to iteratively solve a candidate particle velocity gradient. In P2G, instead of calculating the stress, a material constraint is solved for all particles independently. For example, for an elastic material with MLS-MPM formulation, the co-rotational constraint-solving step includes
@@ III-D Computational Speed Comparison
We provide an initial performance evaluation when simulating a cubic object falling under gravity with different numbers of particles for the implemented solvers on Nvidia RTX 4090 (compute capability 8.9), with CUDA Toolkit version 12.4.1. For all solvers, each step advances the simulation by $0.02$ second. For MLS-MPM and standard MPM, each step contains 10 integration sub-steps with $\Delta t=0.002$. For PB-MPM, each step uses $\Delta t=0.02$ with 10 or 20 iterations. The grid size is fixed in all situations with $175,616$ nodes. The results are shown in Fig. 3. While comparing with existing implementations is reasonable, this work does not aim to outperform existing code in computational speed, as our main focus is on including multiple MPM solver options within one library. Further code optimization and performance benchmarking are planned as future work.
@@ IV-A Soft-Rigid Coupling
Figure 4: Grid level collision with SDF. Nodal velocities are modified if they are inside a rigid body.
Two-way coupling between MPM soft bodies and externally-solved rigid bodies can be achieved at the grid level [29, 33]. The most straightforward approach is to use a signed distance field (SDF) for each rigid body and modify the predicted nodal velocity based on the contact model if a grid node is inside the rigid geometry. This is similar to a predictor-corrector method [10], as shown in Fig. 4. We set the nodal velocity to zero in the normal direction, and apply frictional correction to the tangential component, unless the node is separating from the rigid body. Mathematically, the relative velocity between a node and a rigid body is $\mathbf{v}_{rel}=\mathbf{v}_{I}-\mathbf{v}_{r}$, with $\mathbf{v}_{r}$ being the rigid body’s velocity. Given a rigid surface normal $\mathbf{n}$, $\mathbf{v}_{rel}$ is decomposed into the tangent component $\mathbf{v}_{tg}$ and the normal part $v_{n}$:
@@
Figure 5: Soft body (blue) coupling with rigid bodies (white sphere and cube) of different masses.
@@ IV-B Geometries and Contact Method for Cutting
Additionally, since a knife or scalpel can cut with only the edge side instead of the spine (blunt) side, it is inappropriate to have a geometry that cuts in all directions. Considering this, we apply a circular-shaped distance field to the spine side and use sticky contact (zero relative velocity) if a node is in this region. When the spine side touches an MPM soft body, it generates contact similar to a cylinder colliding with the soft body, as shown in Fig. 6. The SDF of a scalpel’s cross-section is illustrated in Fig. 7. The geometry is named slicer in our code implementation, which includes QuadSlicer and TriangleMeshSlicer.
@@
Figure 7: SDF for a slicer geometry. The signs of the distance on both sides are different.
@@ IV-C Geometries and Contact Method for Suturing
In this work, we assume that the suture thread can be discretized into connected straight-line segments. To achieve suturing, we introduce the arc and the line segment geometries, which can model an arc-shaped suture needle and a discrete thread segment. They are both 1D geometries with SDFs. The signed distance is calculated analytically from a given query point to its nearest projected point on the geometry, which is always positive. The surface normal $\mathbf{n}$ is the normalized vector from the projected point to the query point. Nodal velocities inside the SDF region are corrected differently than for regular geometries. The relative velocity is decomposed into three parts: one along the arc or line tangent $\mathbf{v}_{tg1}$, one that lies on the cross-section of the geometry $\mathbf{v}_{tg2}$, and one along the surface normal $\mathbf{n}v_{n}$, as shown in Fig. 8, and only the tangent component is kept. Mathematically, the relative velocity $\mathbf{v}_{rel}$ is decomposed to
@@
where $\mathbf{v}_{tg1}$ is the true tangent velocity along the arc or line. Nodal velocity correction is applied by replacing $\mathbf{v}_{tg}$ with $\mathbf{v}_{tg1}$ in (18). This ensures that the rigid geometry can move inside a soft body if the motion is along the geometry (e.g., inserting a needle) while preventing lateral motion, as shown in Fig. 9. Geometries for suturing include Arc and ConnectedLineSegments, where the latter is a generalized geometry of a line segment.
@@ V-A Object Model and Simulation Lifecycle
The library follows an object model similar to the one in PhysX, such that integration with other PhysX-based software is straightforward, as shown in Fig. 10.
@@ V-B Implementation Details and Usage
The library is implemented in CUDA C++ and is dependent only on the standard C++ library and the CUDA runtime. 3 types of Solver discussed in Section III are currently included. ParticleObject represents an object of material points simulated by MPM, and Shape is a rigid body. SimulationFactory is used to create all objects. C++ Users should only include the header simulation_factory.h. A backend C-style API engine is also provided for integration into existing software.
@@ V-C Unity Integration
Integration into Unity is straightforward using the C-style APIs. Only a few modifications are needed to allow the coupling between the MPM soft body and rigid bodies simulated in Unity’s default engine. For each rigid body in Unity, a corresponding Shape is created for two-way coupling with the soft body. During each step, the rigid body is simulated first with Unity, after which the Shape’s velocity is updated. The MPM engine is then advanced, recording a total momentum change of the contact grids with the rigid body. This is then used to apply a momentum impulse on the rigid body back in Unity. A Unity plugin is provided as part of CRESSim-MPM, as shown in Fig. 11a.